is 225° equal to 7 times pi divided by 6 radians.
Answer by lwsshak3(11628)

is 225° equal to 7 times pi divided by 6 radians
..
7π/6 radians is found by rotating counter-clockwise 6π/6 (180º) plus another π/6 (30º) to get 7π/6 =180+30=210º